OpenStudy (anonymous):

The scale on a house blueprint is 1 inch to 3 feet. How big is the kitchen on the blueprint, if in reality it is 14 ft times 16 ft. pLZZ HELP

OpenStudy (anonymous):

so if theres 1 inch to 3 feet...theres x inches to 14 feet 1/3=x/14 cross multiply 3x=14 x=14/3 or 4 and 2/3 inches now try it with the other side of 16 feet 1/3=x/16 x=?

OpenStudy (anonymous):

wht?